The distribution of actual weights of 8oz. wedges of cheddar cheese produced by a certain company is normal with mean 8.1 ounces and standard deviation 0.1 ounces. If a sample of five of these cheese wedges is selected. What is the probability that their average weight is less than 8oz?
Answer by stanbon(75887)

The distribution of actual weights of 8oz. wedges of cheddar cheese produced by a certain company is normal with mean 8.1 ounces and standard deviation 0.1 ounces. If a sample of five of these cheese wedges is selected. What is the probability that their average weight is less than 8oz?
---------------------------
t(8) = (8-8.1)/[0.1/sqrt(5)] = -2.2361
---
P(x < 8) = P(t < -2.2361 when df = 4) = tcdf(-100,-2.2361,4) = 0.0445
===================
Cheers,
Stan H.
===================